# Deep-ELA: Deep Exploratory Landscape Analysis with Self-Supervised Pretrained Transformers for Single- and Multi-Objective Continuous Optimization Problems

### v1

@@ Abstract
In many recent works, the potential of Exploratory Landscape Analysis (ELA) features to numerically characterize, in particular, single-objective continuous optimization problems has been demonstrated. These numerical features provide the input for all kinds of machine learning tasks on continuous optimization problems, ranging, i.a., from High-level Property Prediction to Automated Algorithm Selection and Automated Algorithm Configuration. Without ELA features, analyzing and understanding the characteristics of single-objective continuous optimization problems would be impossible.
Yet, despite their undisputed usefulness, ELA features suffer from several drawbacks. These include, in particular, (1.) a strong correlation between multiple features, as well as (2.) its very limited applicability to multi-objective continuous optimization problems. As a remedy, recent works proposed deep learning-based approaches as alternatives to ELA. In these works, e.g., point-cloud transformers were used to characterize an optimization problem’s fitness landscape. However, these approaches require a large amount of labeled training data.
@@ 1 Introduction and Related Work
Optimization problems, often found at the heart of numerous scientific and industrial applications, present challenges that necessitate robust and efficient problem-solving methodologies. A central concern in this field is the analysis and characterization of the continuous optimization problem landscapes. One methodological concept that has emerged as particularly significant for such a characterization is Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011). ELA facilitates a numerical representation of single-objective continuous optimization problems, providing insights into their intrinsic characteristics. The numerical features harvested from ELA serve as the foundation for various machine-learning tasks pertinent to continuous optimization. Such tasks include, but are not limited to, High-level Property Prediction (HL2P; Mersmann et al.,, 2011; Seiler et al.,, 2022; Volz et al.,, 2023), Automated Algorithm Selection (AAS; Rice,, 1976; Kerschke et al.,, 2019), and Automated Algorithm Configuration (AAC; Hutter et al.,, 2009; Huang et al.,, 2019; Schede et al.,, 2022).
@@
Given the strengths and weaknesses of both, ELA and deep learning-based methodologies, there is a compelling case to be made for a synthesis of the two. This paper seeks to bridge this gap by introducing a hybrid approach, capitalizing on the merits of both paradigms. We thus present a novel method that employs four pre-trained transformer models tailored to extract deep representations of the landscapes of both single- and multi-objective continuous optimization problems. This innovative approach promises enhanced flexibility and efficacy in the realm of optimization problem analysis and understanding.
We introduce these pre-trained transformers as Deep Exploratory Landscape Analysis (Deep-ELA). The strengths of Deep-ELA are prominently seen in its inherent adaptability to multi-objective continuous optimization problems and the reduced correlation among its learned features. We demonstrate the straightforward adaptability of Deep-ELA by applying it to rigorous tests across three distinct case studies that showcase their efficacy:
@@
The remainder of this paper is structured as follows. Initially, we equip readers with a concise background on ELA, providing an overview of feature-free alternatives and laying out the notation employed throughout the paper (Chapter 2). Subsequently, Chapter 3 dives into the architectural essence of Deep-ELA. Following this, we delve into a thorough exploration of the various datasets featured in this work in Chapter 4. We then pivot to a comparative analysis, contrasting the outcomes of Deep-ELA with traditional ELA across the previously mentioned case studies in Chapter 5. In Chapter 6, we present our conclusions and provide opportunities for future research.
@@ 2 Background
In the following, we will briefly outline the notation of this paper. An optimization function is given by:
@@
Here, $\mathbf{x}^{*}$ is the optimal solution for $Z$. The goal of single-objective optimization is to find a solution $\mathbf{x}^{*}$ that w.l.o.g. minimizes the function $Z$ globally. If multiple optimal solutions exist, $Z$ is multimodal, and the aim shifts to locating all optimal solutions.
@@
In many studies concerning algorithm behavior and problem understanding, $Z$ is treated as a blackbox optimization problem, implying that direct formal analyses cannot be conducted on $Z$. Instead, preliminary analyses are often carried out on a small sample of candidate solutions $X$, which is usually derived using a random or quasi-random generator like Latin Hypercube Sampling (LHS; McKay et al.,, 1979) or Sobol sequences (Sobol’,, 1967). The corresponding image $Y$ is obtained by evaluating the optimization problem $Z$ at the candidate solutions $X$. The tuple $(X,Y)$ is then available for subsequent studies.
@@ 2.1 Exploratory Landscape Analysis
For downstream tasks involving continuous optimization problems, traditional machine learning methods aim to learn a specific underlying task. In AAS, for instance, classical machine learners like Support Vector Machines (SVM; Cortes and Vapnik,, 1995), Random Forests (RF; Breiman,, 2001), or $k$-Nearest Neighbors ($k$NN; Cover and Hart,, 1967), are utilized to address the Algorithm Selection Problem (ASP) (see Figure 1a). However, these classical learners cannot directly process raw samples of candidate solutions, i.e., $(X,Y)$, as the learners necessitate feature vectors as input per problem instance. Yet, transforming the set of (randomly distributed) points into a vector is not suitable, as such a mapping either (a) lacks an unambiguous order of the points, or (b) requires the points to be aligned in a grid structure, which in turn is affected by the curse of dimensionality. Therefore, it is essential to convert the set of candidate solutions $(X,Y)$ into a vector of numerical values that provides a meaningful representation of the candidate tuple. This conversion, or transformation, is performed by a mapping function, or short mapper, $F:(\mathcal{X},\mathcal{Y})\to\mathcal{F}\subseteq\mathbb{R}^{{n_{\mathcal{F}}}}$, which translates the decision and objective values into a fixed-length real-valued vector, termed (instance) features, in the feature space $\mathcal{F}$.
For continuous optimization problems, Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011) features are commonly employed in many downstream tasks, including AAS. Formally, ELA is a function that maps samples $(X,Y)$ from the decision and objective space $(\mathcal{X},\mathcal{Y})$ to the ${c_{\text{ELA}}}$-dimensional feature space $\mathcal{F}_{\text{ELA}}\subseteq\mathbb{R}^{{c_{\text{ELA}}}}$
@@
Refer to caption (a) ELA Features
Refer to caption (b) Deep-ELA Features
@@
While the utility of ELA features is widely recognized, they also have some limitations (Kerschke and Preuss,, 2023). As, e.g., discussed by Renau et al., (2019) and Eftimov et al., (2020), many ELA features exhibit strong correlations to each other, as illustrated in Figure 2a, or have a low Signal to Noise Ratio (SNR), see Figure 3. Moreover, some features from sets like Classical ELA, Information Content, and PCA are sensitive to random scaling and shifting (Renau et al.,, 2020); see Prager and Trautmann, (2023) for an in-depth analysis of ELA feature sensitivity. These issues necessitate feature selection for many downstream tasks to eliminate noise or redundancy. Finally, ELA features are primarily tailored for single-objective optimization problems, not multi-objective ones. In fact, Kerschke and Trautmann, (2016) used ELA features to characterize bi-objective problems; however, they only considered simple test problems from DTLZ (Deb et al.,, 2005) and ZDT (Zitzler et al.,, 2000), and applied ELA to the single-objective components of these problems, neglecting any interaction between the objectives.
@@ 2.2 Deep Learning-Based Approaches as Alternative to ELA
Figure 3: Signal to Noise Ratio (SNR) of ELA features (left boxplot) on BBOB compared to the SNR values of features from four Deep-ELA models. We used $\text{SNR}=\mu^{2}/\sigma^{2}$ with mean $\mu$ and standard deviation $\sigma$. For $\sigma\simeq 0$, values are imputed with $10^{12}$, which is the highest observed value. Higher values indicate lower noise. SNR values are calculated per feature based on instances $1$ to $20$ and then mean-aggregated over the $24$ functions and four dimensions ($2,3,5,10$).
@@
Compared to deep learning-based feature-free techniques, ELA-based methods usually require significantly fewer training instances under the assumption that these smaller benchmark sets are nonetheless representative and comprehensive. This becomes particularly advantageous when optimization problems are scarce or crafting a training dataset is computationally expensive. While data augmentation can amplify the training set, deep learning methods may still generalize poorly in case of limited training data. Note that in this context, small means a few hundred and large implies tens or even hundreds of thousands of instances.
@@ 3 Deep Exploratory Landscape Analysis
In the following, we aim to merge the advantages of both feature-based and feature-free analysis in the context of continuous optimization. We introduce a large, pre-trained deep learning model trained on millions of randomly generated continuous, single- and multi-objective optimization instances. The generator for these tasks draws inspiration from Tian et al., (2020) and van Stein et al., (2023), but with adaptations to suit our requirements. Our model’s general goal is the automated extraction of instance features from initial samples of the given optimization instance. We dub this methodology Deep Exploratory Landscape Analysis (Deep-ELA).
@@
Deep-ELA is trained to be invariant against scaling, shifts, and rotations, yielding higher SNR values and minimally correlated features (see Figure 2b and Figure 3) with high importance. Hence, feature selection becomes less important – though still potentially valuable in some cases – as will be demonstrated in the experiments (see Chapter 5). To achieve this, a self-supervised learning task is employed, which educates the model to formulate a representative and unique feature vector for an optimization instance – a process often termed feature-learning. Importantly, Deep-ELA is not restricted to single-objective optimization problems and can naturally be applied to multi-objective optimization problems. We will showcase Deep-ELA’s competitive edge against the traditional ELA but also feature-free approaches in existing studies (e.g., Seiler et al.,, 2022; Prager et al.,, 2022). We also applied our model to a bi-objective case analogous to Rook et al., (2022). See Chapter 5 for an outline of all the case studies.
@@
To the best of our knowledge, only one alternative approach, termed DoE2Vec (van Stein et al.,, 2023), parallels our proposed method. However, DoE2Vec varies substantially from Deep-ELA: its model, rooted in a basic autoencoder design, solely focuses on the objective space $\mathcal{Y}$, completely disregarding the decision space $\mathcal{X}$. This reduced informational scope results in a markedly inferior performance compared to the approaches by Seiler et al., (2022). Additionally, the proposed method is not invariant to the order of objective values in $\mathcal{Y}$, potentially affecting its efficacy.
@@
In terms of Deep-ELA’s architecture (refer to Figure 4), we predominantly adhered to the traditional transformer encoder blueprint as suggested by Vaswani et al., (2017). Contrary to the PCT design by Guo et al., (2021) and Seiler et al., (2022), we have embedded extra feed-forward modules to align more closely with the original transformer. Subsequently, we adopted pre-normalization as recommended by Nguyen and Salazar, (2019) and used Gated Linear Units (GLU; Dauphin et al.,, 2017) as the activation function between all feed-forward modules to reduce the training time. For the feature extractor (the last linear layer), we opted for the Tanh activation, mapping all activations to $[-1,1]$ and, hence, dispensing the need for feature normalization in later tasks.
@@
During the training phase, $X$ is drawn randomly uniformly within the bounded decision space $\mathcal{X}$. We favored uniform sampling over quasi-random sampling strategies to (1.) reduce additional computational overhead during the training procedure, (2.) reduce Deep-ELA’s dependence on a particular sampling method, and (3.) enhance the challenge of the self-supervised training task (which is described in more detail in Chapter 3.1.2). After pre-training, more advanced sampling methods, such as LHS and Sobol sampling, can be employed to improve coverage of the decision space.
Deep-ELA can accommodate any sample length of $(X,Y)$ but has a constraint regarding the sum of the search and objective space dimensions: $d+m\leq\nu$. Here, $\nu$ is an additional hyperparameter of Deep-ELA, termed the degree of dimensionality. It specifies the maximum combined dimensionality of $\mathcal{X}$ (denoted by $d$) and $\mathcal{Y}$ (denoted by $m$). Deep-ELA is versatile and can handle varying dimensionalities $d\in[1,\nu]$ and $m\in[1,\nu]$, provided their combined dimensionality does not surpass $\nu$. This flexibility extends to multi-objective optimization problems since the model can accept $m\geq 1$. It is worth noting that the model was trained on problem instances with $d\geq 2$ and $m\geq 1$.

### v2

@@ Abstract
In many recent works, the potential of Exploratory Landscape Analysis (ELA) features to numerically characterize, in particular, single-objective continuous optimization problems has been demonstrated. These numerical features provide the input for all kinds of machine learning tasks on continuous optimization problems, ranging, i.a., from High-level Property Prediction to Automated Algorithm Selection and Automated Algorithm Configuration. Without ELA features, analyzing and understanding the characteristics of single-objective continuous optimization problems is — to the best of our knowledge — very limited.
Yet, despite their usefulness, as demonstrated in several past works, ELA features suffer from several drawbacks. These include, in particular, (1.) a strong correlation between multiple features, as well as (2.) its very limited applicability to multi-objective continuous optimization problems. As a remedy, recent works proposed deep learning-based approaches as alternatives to ELA. In these works, e.g., point-cloud transformers were used to characterize an optimization problem’s fitness landscape. However, these approaches require a large amount of labeled training data.
@@ 1 Introduction and Related Work
Optimization problems, often found at the heart of numerous scientific and industrial applications, present challenges that necessitate robust and efficient problem-solving methodologies. A central concern in this field is the analysis and characterization of optimization problems’ landscapes, i.e. continuous single objective problems but also i.a. combinatorial optimization, mixed-integer or multi-objective optimization. One methodological concept that has emerged as particularly significant for such a characterization is Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011). ELA facilitates a numerical representation of single-objective continuous optimization problems, providing insights into their intrinsic characteristics. The numerical features harvested from ELA serve as the foundation for various machine-learning tasks pertinent to continuous optimization. Such tasks include, but are not limited to, High-level Property Prediction (HL2P; Mersmann et al.,, 2011; Seiler et al.,, 2022; Volz et al.,, 2023), Automated Algorithm Selection (AAS; Rice,, 1976; Kerschke et al.,, 2019), and Automated Algorithm Configuration (AAC; Hutter et al.,, 2009; Huang et al.,, 2019; Schede et al.,, 2022).
@@
Given the strengths and weaknesses of both, ELA and deep learning-based methodologies, there is a compelling case to be made for a synthesis of the two. This paper seeks to bridge this gap by introducing a hybrid approach, capitalizing on the merits of both paradigms. We thus present a novel method that employs four pre-trained transformer models tailored to extract deep representations of the landscapes of both single- and multi-objective continuous optimization problems. This innovative approach promises enhanced flexibility and efficacy of optimization problem analysis and understanding as these models can be used out-of-the-box without the need for additional feature selection or normalization. In addition, learned features do not require expertise to design meaningful feature sets. Instead, one can simply fine-tune Deep ELA on novel optimization problems that are not meaningfully represented by classical ELA .
We introduce these pre-trained transformers as Deep Exploratory Landscape Analysis (Deep-ELA). [^1] An extension of Deep-ELA to the pflacco package can be found here: https://github.com/mvseiler/deep_ela.git. The strengths of Deep-ELA are prominently seen in its inherent adaptability to multi-objective continuous optimization problems and the reduced correlation among its learned features. We demonstrate the straightforward adaptability of Deep-ELA by applying it to rigorous tests across three distinct case studies that showcase their efficacy:
@@
The remainder of this paper is structured as follows. Initially, we equip readers with a concise background on ELA, providing an overview of feature-free alternatives and laying out the notation employed throughout the paper (Section 2). Subsequently, Section 3 dives into the architectural essence of Deep-ELA. Following this, we delve into a thorough exploration of the various datasets featured in this work in Section 4. We then pivot to a comparative analysis, contrasting the outcomes of Deep-ELA with traditional ELA across the previously mentioned case studies in Section 5. In Section 6, we present our conclusions and provide opportunities for future research.
@@ 2 Background
In the following, we will briefly outline the notation of this paper. An optimization function, $Z$, is given by:
@@
Here, $\mathbf{x}^{*}$ is an optimal solution for $Z$. The goal of single-objective optimization is to find a solution $\mathbf{x}^{*}$ that w.l.o.g. minimizes the function $Z$ globally. If multiple optimal solutions exist, $Z$ is multimodal, and the aim shifts to locating all optimal solutions.
@@
In many studies concerning algorithm behavior and problem understanding, $Z$ is treated as a blackbox optimization problem, implying that there is no algebraic expression of $Z$ given. Hence, no formal algebraic analysis can be conducted directly on $Z$. Instead, preliminary analyses are often carried out on a small sample of candidate solutions $X$, which is usually derived using a random or quasi-random generator like Latin Hypercube Sampling (LHS; McKay et al.,, 1979) or Sobol sequences (Sobol’,, 1967). The corresponding image $Y$ is obtained by evaluating the optimization problem $Z$ at the candidate solutions $X$. The tuple $(X,Y)$ is then available for subsequent studies.
@@ 2.1 Exploratory Landscape Analysis
For downstream tasks involving continuous optimization problems, traditional machine learning methods aim to learn a specific underlying task. In AAS, for instance, classical machine learners like Support Vector Machines (SVM; Cortes and Vapnik,, 1995), Random Forests (RF; Breiman,, 2001), or $k$-Nearest Neighbors ($k$NN; Cover and Hart,, 1967), are utilized to address the Algorithm Selection Problem (ASP) (see Figure 1a). The difference between AAS and ASP is that while the latter refers to the challenge of choosing the best algorithm from a set of algorithms for a specific instance, based on performance metrics like speed or solution quality; the former addresses the use of machine learning to automatically select the best algorithm, typically based on problem features. However, these classical learners cannot directly process raw samples of candidate solutions, i.e., $(X,Y)$, as the learners necessitate feature vectors as input per problem instance. Yet, transforming the set of (randomly distributed) points into a vector is not suitable, as such a mapping either (a) lacks an unambiguous order of the points, or (b) requires the points to be aligned in a grid structure, which in turn is affected by the curse of dimensionality. Therefore, it is essential to convert the set of candidate solutions $(X,Y)$ into a vector of numerical values that provides a meaningful representation of the candidate tuple. This conversion, or transformation, is performed by a mapping function, or short mapper, $F:(\mathcal{X},\mathcal{Y})\to\mathcal{F}\subseteq\mathbb{R}^{{n_{\mathcal{F}}}}$, which translates the decision and objective values into a fixed-length real-valued vector, termed (instance) features, in the feature space $\mathcal{F}$.
For continuous optimization problems, Exploratory Landscape Analysis (ELA; Mersmann et al.,, 2011) features are commonly employed in many downstream tasks, including AAS. ELA was termed by Mersmann et al., (2011) but some feature sets that nowadays are considered as subsets of ELA were already introduced before, i.e. Dispersion (Lunacek and Whitley,, 2006) and Fitness Distance Correlation (Jones et al.,, 1995). Formally, ELA is a function that maps samples $(X,Y)$ from the decision and objective space $(\mathcal{X},\mathcal{Y})$ to the ${c_{\text{ELA}}}$-dimensional feature space $\mathcal{F}_{\text{ELA}}\subseteq\mathbb{R}^{{c_{\text{ELA}}}}$
@@
(a) ELA Features
(b) Deep-ELA Features
@@
Though ELA features are frequently used, they also have some limitations (Kerschke and Preuss,, 2023). As, e.g., discussed by Renau et al., (2019) and Eftimov et al., (2020), many ELA features exhibit strong correlations to each other, as illustrated in Figure 2a, or have a low Signal to Noise Ratio (SNR), see Figure 3. Moreover, some features from sets like Classical ELA, Information Content, and PCA are sensitive to random scaling and shifting (Renau et al.,, 2020); see Prager and Trautmann, (2023) for an in-depth analysis of ELA feature sensitivity. These issues necessitate feature selection for many downstream tasks to eliminate noise or redundancy. Finally, ELA features are primarily tailored for single-objective optimization problems, not multi-objective ones. In fact, Kerschke and Trautmann, (2016) used ELA features to characterize bi-objective problems; however, they only considered simple test problems from DTLZ (Deb et al.,, 2005) and ZDT (Zitzler et al.,, 2000), and applied ELA to the single-objective components of these problems, neglecting any interaction between the objectives.
Another use-case for ELA features is Instance Space Analyzis (ISA; Smith-Miles et al.,, 2014). It can be used to analyze the space that is covered by all instances within a given dataset to identify (dis-)similarities of structural properties between instances. Smith-Miles and Muñoz, (2023) propose a framework that utilizes ELA features for ISA. On the other hand, ISA can also be used to identify empty regions that are not covered by existing optimization problems. (Prager et al.,, 2023) proposed a method to create artificial benchmark problems that cover specific regions within the instance space. These problem instances are learned by neural networks.
@@ 2.2 Deep Learning-Based Approaches as Alternative to ELA
Figure 3: Signal to Noise Ratio (SNR) of ELA features (left boxplot) on BBOB compared to the SNR values of features from four Deep-ELA models. We used $\text{SNR}=\mu^{2}/\sigma^{2}$ with mean $\mu$ and standard deviation $\sigma$. For $\sigma\simeq 0$, values are imputed with $10^{12}$, which is the highest observed value. Higher values indicate lower noise. SNR values are calculated per feature based on instances $1$ to $20$ and then mean-aggregated over the $24$ functions and four dimensions ($2,3,5,10$). Notches show the 95% confidence intervals around the median. The large models yield the highest SNR while the medium models yield the lowest which is to be expected as the large models contain more parameters to create more sophisticated features. Classical ELA features are somewhat ‘in-between’ while simultaneously containing features with multiple, very low SNR values.
@@
Compared to deep learning-based feature-free techniques, ELA-based methods usually require significantly fewer training instances under the assumption that these smaller benchmark sets are nonetheless representative and comprehensive. This becomes particularly advantageous when optimization problems are scarce or crafting a training dataset is computationally expensive. While data augmentation can amplify the training set, deep learning methods may still generalize poorly in case of limited training data. Note that in this context, small means a few hundred and large implies tens or even hundreds of thousands of instances; e.g. Seiler et al., (2022) utilized 144 000144000144\,000144 000 training examples while Prager et al., (2022) only relied on 4 800 training examples (both including ten repetitions per instance).
@@ 3 Deep Exploratory Landscape Analysis
In the following, we aim to merge the advantages of both feature-based and feature-free analysis in the context of continuous optimization. We introduce a large, pre-trained deep learning model trained on millions of randomly generated continuous, single- and multi-objective optimization instances. The generator for these tasks draws inspiration from Tian et al., (2020) and van Stein et al., (2023), but with adaptations to suit our requirements. Our model’s general goal is the automated extraction of instance features from initial samples of the given optimization instance. We dub this methodology Deep Exploratory Landscape Analysis (Deep-ELA). Seiler et al., (2024) analyzed these learned features in detail and, especially, analyzed the complementarity and synergies between classical and Deep ELA.
3.1 Outline of Deep ELA
@@
Deep-ELA is trained to be invariant against scaling, shifts, and rotations, yielding higher SNR values and minimally correlated features (see Figure 2b and Figure 3) with high importance. Hence, feature selection becomes less important — though still potentially valuable in some cases, see Seiler et al., (2024) — as will be demonstrated in the experiments (see Section 5). To achieve this, a self-supervised learning task is employed, which educates the model to formulate a representative and unique feature vector for an optimization instance – a process often termed feature-learning. Importantly, Deep-ELA is not restricted to single-objective optimization problems and can naturally be applied to multi-objective optimization problems. We will showcase Deep-ELA’s competitive edge against the traditional ELA but also feature-free approaches in existing studies (e.g., Seiler et al.,, 2022; Prager et al.,, 2022). We also applied our model to a bi-objective case analogous to Rook et al., (2022). See Section 5 for an outline of all the case studies.
@@
To the best of our knowledge, only two alternative approaches, termed DoE2Vec (van Stein et al.,, 2023) and TransOpt (Cenikj et al.,, 2023), parallels our proposed method. However, DoE2Vec varies substantially from Deep-ELA: its model, rooted in a basic autoencoder design, solely focuses on the objective space $\mathcal{Y}$, completely disregarding the decision space $\mathcal{X}$. This reduced informational scope results in a markedly inferior performance compared to the approaches by Seiler et al., (2022). Additionally, the proposed method is not invariant to the order of objective values in $\mathcal{Y}$, potentially affecting its efficacy. Further, Cenikj et al., (2023) propose also Transformer models, TransOpt, that were trained on an AAS task but the learned (hidden) features can also be used as supplementary for classical ELA features. Contrary to Deep-ELA, the learned features are specific to the underlying AAS task and may lack generalization to other downstream tasks.
@@
In terms of Deep-ELA’s architecture (refer to Figure 4), we predominantly adhered to the traditional transformer encoder blueprint as suggested by Vaswani et al., (2017). Contrary to the PCT design by Guo et al., (2021) and Seiler et al., (2022), we have embedded extra feed-forward modules to align more closely with the original transformer. According to a recent study (Geva et al.,, 2020), the feed-forward layers in transformers act like a memory that stores learned patterns; lower layers store more basic patterns while higher layers store more complex ones. In our case, we expect Deep-ELA to learn and memorize certain patterns of optimization problems. Subsequently, we adopted pre-normalization as recommended by Nguyen and Salazar, (2019) and used Gated Linear Units (GLU; Dauphin et al.,, 2017) as the activation function between all feed-forward modules to reduce the training time as demonstrated by the authors. For the feature extractor (the last linear layer), we opted for the Tanh activation, mapping all activations to $[-1,1]$ and, hence, dispensing the need for feature normalization in later tasks.
@@
During the training phase, $X$ is drawn randomly uniformly within the bounded decision space $\mathcal{X}$. We favored uniform sampling over quasi-random sampling strategies to (1.) reduce additional computational overhead during the training procedure, (2.) reduce Deep-ELA’s dependence on a particular sampling method, and (3.) enhance the challenge of the self-supervised training task (which is described in more detail in Section 3.2.2). After pre-training, more advanced sampling methods, such as LHS and Sobol sampling, can be employed to improve coverage of the decision space.
Deep-ELA can accommodate any sample length of $(X,Y)$ but has a constraint regarding the sum of the decision and objective space dimensions: $d+m\leq\nu$. Here, $\nu$ is an additional hyperparameter of Deep-ELA, termed the degree of dimensionality. It specifies the maximum combined dimensionality of $\mathcal{X}$ (denoted by $d$) and $\mathcal{Y}$ (denoted by $m$). Deep-ELA is versatile and can handle varying dimensionalities $d\in[1,\nu]$ and $m\in[1,\nu]$, provided their combined dimensionality does not surpass $\nu$. This flexibility extends to multi-objective optimization problems since the model can accept $m\geq 1$. It is worth noting that the model was trained on problem instances with $d\geq 2$ and $m\geq 1$.